OpenStudy (anonymous):

Verify 2cos2x/sin2x=cotx-tanx

OpenStudy (anonymous):

is it cos^2x or cos 2x.

OpenStudy (anonymous):

cos2x. no exponent just number

OpenStudy (anonymous):

ok. so cos 2x = 2 cos^2x - 1 which is 2 cos^2x - (sin^2x+cos^2x) Hence cos 2x = cos^2x - sin^2x. also, sin 2x = 2 sin x cos x. So we can simplify the fraction as - \[\frac{ 2(\cos^2x-\sin^2x) }{ 2 \sin x \cos x } => \frac{ (\cos^2x-\sin^2x }{ \sin x \cos x }\]

OpenStudy (anonymous):

ok till here?

OpenStudy (anonymous):

yes! thank you!

OpenStudy (anonymous):

Now you can separate the fractions - \[\frac{ \cos^2x }{ \sin x \cos x } - \frac{ \sin^2x }{ \sin x \cos x }\]

OpenStudy (anonymous):

cancel the cos in LH and sin in RH. you'll get your result. :)

OpenStudy (anonymous):

thank you so much